OpenStudy (mathmath333):

A vessel is full of a mixture of kerosene and petrol in which there is 18% kerosene.Eight litres are drawn off and then the vessel is filled with petrol.If the kerosene is now 15% how much does the vessel hold? a)40 lit b)32 lit c)36lit d)48lit

OpenStudy (mathmath333):

initially kerosene=18% so petrol =82% after kerosene =15% petrol=85% so 3% =8 lit so total kerosene in the mixture =(8x18/3) =48 lit

OpenStudy (mathmath333):

somebody please explain the above solution

OpenStudy (mathmath333):

@amistre64

OpenStudy (amistre64):

from yesterday k+p = initial amount k/(k+p) = .18 p/(k+p) = .82 (k+p) - 8 + Lp is the new amount k/(k+p) - 8 + Lp = .15 p/(k+p) - 8 + Lp = .85

OpenStudy (amistre64):

forogt a set ofo () ...

OpenStudy (amistre64):

k/.18 = p/.82 k = 9p/41 (9p/41)/((9p/41+p) - 8 + Lp) = .15 p/((9p/41+p) - 8 + Lp) = .85 hmm, the wolf says no solution exists

OpenStudy (amistre64):

what is the error in my process?

OpenStudy (mathmath333):

i think when 8 litres of kerosene is taken ,then 8 litres of petrol is filled

OpenStudy (amistre64):

hmm, i cant see that in the instructions, but its not a bad assumption

OpenStudy (amistre64):

removing 8L doesnt change our percentage, only adding in the new p does

OpenStudy (mathmath333):

can u explain the solution given in first comment,

OpenStudy (amistre64):

how can we say that 3% = 8L?

OpenStudy (mathmath333):

because when 8 litres is taken kerosene $ reduces from 18 to 15

OpenStudy (amistre64):

also, the question seems to ask how much does the vessel hold, and not how much kerosene is in the vessel. the amount of kerosene hasnt changed any

OpenStudy (amistre64):

if we have k and p in the same container, and remove 8L we havent changed the ratio by removing some of the contents

OpenStudy (amistre64):

its only in adding new p that the ratio gets changed

OpenStudy (mathmath333):

in the 8 litres removed there can be petrol and kerosene but petrol is filled after that and kersosene % is now 15 so it can be 8 L= 3 %

OpenStudy (amistre64):

if we start with 100L that has a ratio of 18k to 82p then removing 8L still leaves us with 92L with a ratio of 18k to 82p

OpenStudy (amistre64):

adding p to this mixture will give us 92 + L, at a ratio of 15k to 85p

OpenStudy (amistre64):

removing 8L means that we removed: 8(.18)k and 8(.82)p the ratio remains the same, but the quantity reduces

OpenStudy (amistre64):

k/(k+p) = .18 p/(k+p) = .82 k/(k-8(.18)k + p-8(.82)p + Lp) = .15 p/(k-8(.18)k + p-8(.82)p + Lp) = .85

OpenStudy (mathmath333):

i found this equation somewhere 18[1-(8/x)] = 15

OpenStudy (mathmath333):

also found this but i cant understand Kerosene initial = 0.18 of Capacity of Vessel... When 8 litres are withdrawn,Total Kerosene withdrawn=0.18*8=1.44 litres and total Petrol withdrawn = 8-1.44 = 6.56 litres.. After this percentage of kerosene = .15 of Capacity of Vessel... Thus 1.44 litres --> 3 % Hence 1% --> .48 litres hence 100% --> 48 litres..

OpenStudy (amistre64):

i can only explain my own ideas .. lets try this new setup with k=9p/41 since we have reduced the original amount properly while retaining the same initial construction

OpenStudy (amistre64):

wolf hates me ... ugh. im at a loss for it and im not going to be able to explain someone elses approach.

OpenStudy (amistre64):

the first part seems to be the same conclusion i came to, removing 8 removes proportionate parts

OpenStudy (amistre64):

i dont follow the reasoning afterwards tho .. somewhat, but it seems to be making an assumption to me is all

OpenStudy (amistre64):

1.44L of k was removed. but not sure why that is 3% of the new mixture

OpenStudy (amistre64):